DETERMINE THE MOLARITY OF A SOLUTION CONTAINING 2.40 G OF SODIUM CHLORIDE IN 40.0 ML OF SOLUTION.


1
Expert's answer
2022-03-27T10:31:42-0400

M(NaCl)=23+35.5=58.5 g/mol

n(NaCl)=2.4 /58.5=0.04 mol

40 ml=0.04 l

c(NaCl)=0. 04/0.04=1 mol/l